# Witnessing mass-energy equivalence with trapped atom interferometers

### v2

@@ Appendix C Phases
over the ground state trajectory on the corresponding height. For the free-fall trajectory starting at $t_{\text{A}}$ and ending at $t_{\text{B}}$ the correction $\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})$ to the propagation phase reads (compare with (20))
$\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=-\frac{\delta m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right].$ (40)
For a full single oscillation in the optical lattice, the propagation phase correction $\delta_{\text{p,1}}$ is given by
$\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (41)
where $z_{\text{B}}$ and $v_{\text{B}}$ are the atom’s position and velocity at the interaction point, and $\tau_{\text{B}}$ is the period of the oscillation. We assume that the total number $N$ of Bloch oscillations is large, and neglect the correction coming from the fact that the total time $T_{\text{B}}$ of Bloch oscillations is not necessarily equal to the integer multiple of Bloch periods (instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$; however, since $\tau\ll N\tau_{\text{B}}$, we will assume $\tau\approx 0$). Then, the total propagation phase correction is given by
$\delta_{\text{p,tot}}=N\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)N\tau_{\text{B}}\approx\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)T_{\text{B}}.$ (42)
Without loss of generality, we can assume that the interaction height corresponding to the lower trajectory $z_{\text{B}}^{\text{(d)}}=0$, while the interaction height of the upper one equals $z_{\text{B}}^{\text{(u)}}=\Delta z$. We introduce the mean velocity (within a single oscillation) $\langle v^{2}\rangle\equiv v_{\text{B}}^{2}/3$, and note that $\delta m=\omega_{0}/c^{2}$. Combining the propagation phase with the laser phase corrections, we get
$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}}.\end{split}$ (43)
@@ Appendix D Non-perturbative calculations
Since the phase acquired due to the interaction with the clock pulses is position-independent, it will not change in the non-perturbative analysis. Therefore, we consider only the phases acquired during the Bloch oscillations within the optical lattice. The trajectory followed by both trajectories has been described above. We analyze only the time between the two first interactions on the ground state trajectory (at some arbitrary height $z_{\text{B}}$) and show that the phase difference between the ground and the excited state trajectories agrees with the perturbative calculations. It is straightforward that it will agree with the perturbative calculations also at later times.
@@
(this is just Eq. (20) with $t_{\text{A}}\to 0$ and $t_{\text{B}}\to t$) where $m^{\text{(g)}}=m$ and $m^{\text{(e)}}=m+\delta m$ are the masses of the atom in the ground and excited state, respectively, $v_{0}$ is the initial velocity, and $z_{0}$ is the initial height. For the ground state, we replace $v_{0}$ by $v_{\text{B}}$, while for the excited state by $v_{\text{B}}-\delta v_{\text{B}}$. The initial height for both states is the same and equal to $z_{\text{B}}$.
@@
$\delta_{\text{p}}(t)=\phi_{\text{p}}^{\text{(e)}}(t)-\phi_{\text{p}}^{\text{(g)}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right]-\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t$ (46)
(recall that we neglect the terms higher-order than $1/c^{2}$). Let us notice that, since $\delta v_{\text{B}}t$ is just the separation $\Delta z(t)$ between the trajectories, the second term in the above formula is equal in value but with the opposite sign to the separation phase
$\delta_{\text{s}}(t)=\frac{\bar{p}(t)\delta v_{\text{B}}t}{\hbar}=\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t,$ (47)
where $\bar{p}(t)=m(v_{\text{B}}-gt)+\mathcal{O}(1/c^{2})$. Therefore, the total phase difference for $t<\tau_{\text{B}}-\delta v_{\text{B}}/g$ is given by
$\delta_{\text{p}}(t)+\delta_{\text{s}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (48)
@@
At the time $\tau_{\text{B}}-\delta v_{\text{B}}/g$ the interaction of the atom on the excited trajectory with the laser occurs, and it acquires a laser phase
$\phi_{\text{l}}^{\text{(e)}}=\frac{m}{\hbar}2v_{\text{B}}(z_{\text{B}}-\delta z_{\text{B}})=\phi_{\text{l}}^{\text{(g)}}-\frac{m}{\hbar}\delta v_{\text{B}}\tau_{\text{B}}+\mathcal{O}(\delta m^{2}),$ (49)
where $\phi_{\text{l}}^{\text{(g)}}$ is the laser phase acquired by the ground state trajectory at time $\tau_{\text{B}}$. Let us notice that right before the second interaction of the excited state the separation phase is equal to
$\delta_{\text{s}}(\tau_{\text{B}}-\Delta v/g-\varepsilon)=-\frac{m}{\hbar}v_{\text{B}}\delta v_{\text{B}}t,$ (50)
while right after the interaction it is $\sim\mathcal{O}(1/c^{4})$, because the mean momentum is then $\sim\mathcal{O}(1/c^{2})$. This situation (mean momentum $\sim\mathcal{O}(1/c^{2})$ and thus negligible value of separation phase) persists until the second interaction on the ground state trajectory, at which time both trajectories are at the same height, but the excited one has the velocity $v_{\text{B}}-2\delta v_{\text{B}}$. At this point, we can repeat the whole analysis changing everywhere $\delta v_{\text{B}}\to 2\delta v_{\text{B}}$. Since the propagation phase difference acquired between the second interaction on the excited and on the ground trajectories is also $\sim\mathcal{O}(\delta m^{2})$, the total phase difference after the first oscillation is equal to
$\delta=\delta_{\text{p}}+\delta_{\text{s}}+\delta_{\text{l}}=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)\tau_{\text{B}}-v_{\text{B}}g\tau_{\text{B}}^{2}+\frac{1}{3}g^{2}\tau_{\text{B}}^{3}\right]=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (51)

### v3

@@ Appendix C Phases
over the ground state trajectory on the corresponding height. For the free-fall trajectory starting at $t_{\text{a}}$ and ending at $t_{\rm B}$ the correction $\delta_{\text{p}}(t_{\text{a}}\to t_{\rm B})$ to the propagation phase reads (compare with (20))
$\delta_{\text{p}}(t_{\text{a}}\to t_{\rm B})=-\frac{\delta m}{\hbar}\left[\left(c^{2}+gz_{\text{a}}-\frac{1}{2}v_{\text{a}}^{2}\right)(t_{\rm B}-t_{\text{a}})+v_{\text{a}}g(t_{\rm B}-t_{\text{a}})^{2}-\frac{1}{3}g^{2}(t_{\rm B}-t_{\text{a}})^{3}\right].$ (40)
For a full single oscillation in the optical lattice starting at height $z_{\rm B}$ with velocity $v_{\rm B}$, the propagation phase correction $\delta_{\text{p,1}}$ is given by
$\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left[-c^{2}-g\left(z_{\rm B}+\frac{v_{\rm B}^{2}}{3}\right)+\frac{v_{\rm B}^{2}}{6}\right]\tau_{\rm B}\equiv-\omega_{0}\tau_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right],$ (41)
where $\tau_{\rm B}$ is the oscillation period, we have used the fact that $\delta m=\hbar\omega_{0}/c^{2}$, and introduced $\langle z\rangle$ and $\langle v^{2}\rangle$ for the mean height and velocity squared. We assume that the total number $N$ of Bloch oscillations is large, and neglect the correction coming from the fact that the total time $T_{\rm B}$ of Bloch oscillations is not necessarily equal to the integer multiple of Bloch periods (instead, $T_{\rm B}=N\tau_{\rm B}+\tau$ with $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$; however, since $\tau\ll N\tau_{\rm B}$, we will assume $\tau\approx 0$). Then, the total propagation phase correction is given by
$\delta_{\text{p,tot}}\approx N\delta_{\text{p,1}}=-\omega_{0}N\tau_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right]\approx-\omega_{0}T_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right].$ (42)
Without loss of generality, we can assume that the mean height corresponding to the lower trajectory $\langle z\rangle_{\rm d}=0$, while this of the upper one equals $\langle z\rangle_{\rm u}=\Delta z$. Combining the propagation phase with the laser phase corrections, we get
$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1+\frac{g\Delta z}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B}.\end{split}$ (43)
@@ Appendix D Non-perturbative calculations
Since the phase acquired due to the interaction with the clock pulses is position-independent, it will not change in the non-perturbative analysis. Therefore, we consider only the phases acquired during the Bloch oscillations within the optical lattice. The trajectory followed by both trajectories has been described above. We analyze only the time between the two first interactions on the ground state trajectory (at some arbitrary height $z_{\rm B}$) and show that the phase difference between the ground and the excited state trajectories agrees with the perturbative calculations. It is straightforward that it will agree with the perturbative calculations also at later times.
@@
(this is just Eq. (20) with $t_{\text{A}}\to 0$ and $t_{\rm B}\to t$) where $m^{\text{(g)}}=m$ and $m^{\text{(e)}}=m+\delta m$ are the masses of the atom in the ground and excited state, respectively, $v_{0}$ is the initial velocity, and $z_{0}$ is the initial height. For the ground state, we replace $v_{0}$ by $v_{\rm B}$, while for the excited state by $v_{\rm B}-\delta v_{\rm B}$. The initial height for both states is the same and equal to $z_{\rm B}$.
@@
$\delta_{\text{p}}(t)=\phi_{\text{p}}^{\text{(e)}}(t)-\phi_{\text{p}}^{\text{(g)}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)t-v_{\rm B}gt^{2}+\frac{1}{3}g^{2}t^{3}\right]-\frac{m}{\hbar}(v_{\rm B}-gt)\delta v_{\rm B}t$ (46)
(recall that we neglect the terms higher-order than $1/c^{2}$). Let us notice that, since $\delta v_{\rm B}t$ is just the separation $\Delta z(t)$ between the trajectories, the second term in the above formula is equal in value but with the opposite sign to the separation phase
$\delta_{\text{s}}(t)=\frac{\bar{p}(t)\delta v_{\rm B}t}{\hbar}=\frac{m}{\hbar}(v_{\rm B}-gt)\delta v_{\rm B}t,$ (47)
where $\bar{p}(t)=m(v_{\rm B}-gt)+\mathcal{O}(1/c^{2})$. Therefore, the total phase difference for $t<\tau_{\rm B}-\delta v_{\rm B}/g$ is given by
$\delta_{\text{p}}(t)+\delta_{\text{s}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)t-v_{\rm B}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (48)
@@
At the time $\tau_{\rm B}-\delta v_{\rm B}/g$ the interaction of the atom on the excited trajectory with the laser occurs, and it acquires a laser phase
$\phi_{\text{l}}^{\text{(e)}}=\frac{m}{\hbar}2v_{\rm B}(z_{\rm B}-\delta z_{\rm B})=\phi_{\text{l}}^{\text{(g)}}-\frac{m}{\hbar}\delta v_{\rm B}\tau_{\rm B}+\mathcal{O}(\delta m^{2}),$ (49)
where $\phi_{\text{l}}^{\text{(g)}}$ is the laser phase acquired by the ground state trajectory at time $\tau_{\rm B}$. Let us notice that right before the second interaction of the excited state the separation phase is equal to
$\delta_{\text{s}}(\tau_{\rm B}-\Delta v/g-\varepsilon)=-\frac{m}{\hbar}v_{\rm B}\delta v_{\rm B}t,$ (50)
while right after the interaction it is $\sim\mathcal{O}(1/c^{4})$, because the mean momentum is then $\sim\mathcal{O}(1/c^{2})$. This situation (mean momentum $\sim\mathcal{O}(1/c^{2})$ and thus negligible value of the separation phase) persists until the second interaction on the ground state trajectory, at which time both trajectories are at the same height, but the excited one has the velocity $v_{\rm B}-2\delta v_{\rm B}$. At this point, we can repeat the entire analysis changing everywhere $\delta v_{\rm B}\to 2\delta v_{\rm B}$. Since the propagation phase difference acquired between the second interaction on the excited and on the ground trajectories is also $\sim\mathcal{O}(\delta m^{2})$, the total phase difference after the first oscillation is equal to
$\delta=\delta_{\text{p}}+\delta_{\text{s}}+\delta_{\text{l}}=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)\tau_{\rm B}-v_{\rm B}g\tau_{\rm B}^{2}+\frac{1}{3}g^{2}\tau_{\rm B}^{3}\right]=\frac{\delta m}{\hbar}\left[-c^{2}-g\left(z_{\rm B}+\frac{v_{\rm B}^{2}}{3}\right)+\frac{v_{\rm B}^{2}}{6}\right]\tau_{\rm B},$ (51)
@@
Appendix E Bloch oscillations — the usual description
In our analysis, we have employed a simplified model of Bloch oscillations based on [61, 19], where the oscillating atom follows a bouncing trajectory in a gravitational field (see Fig. 2). However, this is not the most common description of Bloch oscillations — usually one assumes that the dynamics of the atom is set by a periodic potential (optical lattice) and a linear correction (gravitational field), which leads to the oscillatory motion [68, 69]. The full Hamiltonian of the system involving the leading relativistic corrections reads
[…]
where $V_{0}$ is the depth of the optical lattice potential, $k_{\text{Bloch}}$ is the wave vector of the lattice photons, $|\ket{e}\bra{e}$ is a projector on the excited state of the atom, and
[…]
Writing Eq. (52) we have assumed that the lattice operates at the “magic” wavelength [58, 72], which means that the harmonic potential experienced by the ground and excited state atoms is the same. Let us notice that the frequency of the light as perceived by the atom depends on its position and velocity (due to the relativistic Doppler shifts). Therefore, in principle, it can only be at the “magic” wavelength for an atom at some specific height and velocity. The atom at some other height, or moving with a different velocity, will perceive the light as shifted from this wavelength, leading to a dependence of the optical lattice potential on the center-of-mass degrees of freedom. Note that this is exactly analogous to the situation in optical lattice clocks, where the atoms are not perfectly static and are placed at different heights. However, it was shown in [58] that around the “magic” wavelength the potential changes very slowly with the frequency of the lattice laser, and the corrections due to Doppler shifts can be made negligible. Consequently, we shall omit them from further consideration.
[…]
$\delta_{\rm d}^{(0)}=\delta_{\rm u}^{(0)}=(\omega-\omega_{0})T_{\rm B}.$ (54)
The first-order (in $1/c^{2}$) correction to the phases results from the correction to the eigenenergies and is given by
$\delta_{i}^{(1)}=-\frac{1}{\hbar}\left(\left\langle\hat{H}^{(1)}\right\rangle_{e,i}-\left\langle\hat{H}^{(1)}\right\rangle_{g,i}\right)T_{\rm B},$ (55)
where the expectation values are evaluated on unperturbed states. In the ground state, the expectation value of $\hat{H}^{(1)}$ is zero irrespectively of the path, and in the excited state we have
$\left\langle\hat{H}^{(1)}\right\rangle_{e,i}=\frac{\hbar\omega_{0}}{c^{2}}\left(g\langle\hat{z}\rangle_{i}-\frac{\langle\hat{p}^{2}\rangle_{i}}{2m^{2}}\right).$ (56)
This leads to the total phase shift of the form:
$\delta_{i}=\delta_{i}^{(0)}+\delta_{i}^{(1)}=\left[\omega-\omega_{0}\left(1+\frac{g\langle\hat{z}\rangle_{i}}{c^{2}}-\frac{\langle\hat{p}^{2}\rangle_{i}}{2m^{2}c^{2}}\right)\right]T_{\rm B},$ (57)
which, upon identifications $\langle\hat{z}\rangle_{i}$ and $\langle\hat{p}^{2}\rangle_{i}=m^{2}\langle v^{2}\rangle_{i}$, takes the same form as derived in Appendix C. Note the different meaning of the brackets in (57) and those present e.g. in (43) — in the first case they indicate averaging over the quantum state, while in the latter they mean taking the average over the classical trajectory.
The same results can be derived in the Lagrangian formalism using the method developed in [62]. We assume that the atom follows a superposition of classical trajectories set by the (ground-state) Lagrangian
$L_{g}=-mc^{2}+\frac{1}{2}mv^{2}-mgz-V_{0}\cos(2k_{\text{Bloch}}z),$ (58)
corresponding to the Hamiltonian (52). Let us consider (following the reasoning from Appendix C) the dominant correction to the Lagrangian of the excited state atoms, i.e.,
$\delta L=-\delta m\left(c^{2}+gz-v^{2}/2\right).$ (59)
To calculate the phases $\delta_{\rm d}$ and $\delta_{\rm u}$ ($\Delta\phi$ does not depend on the model of Bloch oscillations that we use) we integrate the Lagrangian correction $\delta L$ over the trajectory $z(t)$ followed by the atom (assuming that the trajectory is the same for both internal states). This gives (recall that $\delta m=\hbar\omega_{0}/c^{2}$)
$\frac{1}{\hbar}\int_{0}^{T_{\rm B}}\mathrm{d}t\delta L=-\omega_{0}T_{\rm B}\left[1+\frac{g}{c^{2}}\left(\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tz(t)\right)-\frac{1}{2c^{2}}\left(\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tv^{2}(t)\right)\right]\equiv-\omega_{0}T_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right],$ (60)
where we have introduced
$\langle z\rangle\equiv\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tz(t),\qquad\langle v^{2}\rangle\equiv\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tv^{2}(t).$ (61)
Assuming that the mean height of the lower trajectory is set to zero and that of the upper one to $\Delta z$, that the mean squared velocity is the same on both heights and adding a laser phase $\omega T_{\rm B}$ coming from the interaction with the clock pulse, we get
$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1+\frac{g\Delta z}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\end{split}$ (62)
which is the same as derived in Appendix C. It is worth noting that the method developed in [62] was derived for quadratic Lagrangians, while Lagrangian (58) is nonquadratic. However, the compatibility with the result obtained in the Hamiltonian formalism suggests that the method from [62] is still applicable in our case.